# Data Readiness for AI: A 360-Degree Survey

### v1

@@ 5. Gaps and Challenges
This section discusses the gaps and challenges in the metrics for assessing data readiness for AI in structured and unstructured data. While structured data poses unique challenges regarding standardization, interpretability, and context sensitivity, unstructured datasets present additional complexities due to their diverse formats, varying modalities, and contextual nuances. By exploring the gaps and challenges in both structured and unstructured data readiness metrics, we aim to discuss the key areas that require further research, collaboration, and innovation in AI and data science.
@@
Assessing data readiness for AI and data science applications, regardless of its structure, presents several significant challenges origins from the absence of a unified framework that complicates the comparison and consistent application of existing metrics across diverse dimensions. This absence hampers the development of a cohesive and comprehensive data readiness assessment method explicitly tailored for different types of data. Although Gupta et al.’s (Gupta et al., 2021) data quality toolkit has taken initial steps in addressing this concern, its coverage is primarily focused on structured data, leaving a need for further advancements to include a broader range of dimensions related to structured data readiness and extend the toolkit’s capabilities to address unstructured data challenges.
@@
In the rapidly evolving fields of AI and data science, emergence of new use cases and diverse data structures constantly challenges the evaluation of data readiness. To ensure these metrics remain effective in assessing data preparedness for the latest AI applications, they must adapt and evolve alongside the technology. Striking the right balance between data quality and quantity is another crucial challenge. While numerous metrics focus on data quality, a comprehensive approach should consider the trade-off between data quality and sufficiency. Sufficient data volumes are essential for meaningful analysis and effective AI model training, making finding the optimal equilibrium between data quality and quantity a challenge. Additionally, clear interpretability of these readiness metrics is essential for stakeholders to grasp their implications on overall data quality and their potential impacts on AI model performance. Enhanced visualization and explanation techniques can significantly improve the practical utility of these metrics, facilitating more informed decision-making processes.
@@
Addressing the challenges in the dynamic field of AI and data readiness is important. Specific AI applications often require customized data readiness metrics due to varied data quality standards, necessitating domain-specific expertise for effective navigation. Simultaneously, addressing subjective judgments and human biases, particularly concerning fairness and privacy, adds another layer of complexity. Developing unbiased and ethical metrics that adapt to various data types and applications requires ongoing research and innovation. Furthermore, the ever-evolving nature of real-world datasets require continuous data readiness assessment throughout an AI system’s lifecycle, extending beyond the initial data preparation phase.
@@
Data access and ownership concerns can impede data readiness evaluation, particularly when datasets are restricted due to privacy and ownership issues. These limitations can delay a comprehensive data readiness assessment, requiring collaborative efforts and agreements between data providers and users to navigate these challenges effectively. Furthermore, the ongoing challenge lies in establishing meaningful thresholds that define acceptable data readiness levels. Given the context-dependent nature of data and the diversity of AI use cases, universal and context-independent threshold values are challenging to define. Clear guidelines regarding data quality thresholds are essential to ensure consistent and effective data preparation practices. Lastly, to compare and evaluate various data readiness metrics, developing well-established benchmark datasets and evaluation protocols becomes crucial. Creating representative benchmarks spanning different industries and data structures can facilitate a more comprehensive comparison of diverse metrics and scoring mechanisms.
@@ 6. Conclusion
This comprehensive survey underscores the critical role of data preparation in the success of AI applications. We explored the challenges, tools, and metrics associated with data readiness, emphasizing its significance in achieving accurate and dependable AI-driven outcomes. Our study highlights the need for a holistic approach, covering structured and unstructured data, and underscores the importance of incorporating fairness-related metrics to ensure unbiased AI decision-making. We have identified and showcased existing metrics and scoring mechanisms that effectively measure data readiness available in literature by studying more than 120 sources from reputable journals including ACM, IEEE, and other reputable journals, as well as web articles, spanning the past three decades. By thoroughly exploring these dimensions and metrics, we have contributed to a deeper understanding of data readiness within AI applications. As AI advances and data becomes an even more critical asset, staying up to date with the latest research and advancements in data readiness metrics is essential. This survey provides a foundational resource for researchers and practitioners, equipping them with the essential insights needed to navigate the complexities of data preparation for AI effectively and emphasizing that data readiness is not just a preliminary step but an ongoing commitment.

### v2

@@ 6. Gaps and Challenges
We discuss the challenges in defining the metrics for assessing data readiness for AI in structured and unstructured data. While structured data poses unique challenges regarding standardization, interpretability, and sensitivity, unstructured datasets present additional complexities due to their diverse formats, varying modalities, and contextual nuances.
@@
Assessing data readiness for AI and data science applications, regardless of its structure, presents several challenges stemming from the absence of a unified framework that complicates the comparison and consistent application of metrics across diverse dimensions. This absence hampers the development of a cohesive and comprehensive data readiness assessment method explicitly tailored for different types of data. Although IBM DQT (et al., 2021) and AIDRIN (Hiniduma et al., 2024) have taken the initial steps in addressing this concern, their coverage is primarily focused on structured data, leaving a need for further advancements to include a broader range of dimensions related to structured data readiness and extend the toolkit’s capabilities to address unstructured data challenges. Additionally, with the rapid growth of large datasets, there is a lack of parallel systems designed to evaluate data readiness effectively at scale. This highlights the need for further advancements to develop robust methods for handling diverse and extensive data environments.
A significant challenge in DRAI assessment is the evolving nature of data quality dimensions. As new data quality metrics emerge, they add complexity to the evaluation process. According to Batini and Scannapieco’s (Batini and Scannapieco, 2006), the development of metrics specific to various domains, such as the condition of archival documents, the integrity of statistical data, and the positional accuracy in geospatial data, further complicates comprehensive assessments of data quality and readiness. This ongoing evolution necessitates continual adaptation and refinement of evaluation frameworks to effectively address established and new metrics.
@@
In the rapidly evolving fields of AI and data science, the emergence of new use cases and diverse data structures constantly challenges the evaluation of data readiness. To ensure these metrics remain effective in assessing data preparedness for the latest AI applications, they must adapt and evolve alongside the technology. Striking the right balance between data readiness and quantity is another crucial challenge. While numerous metrics focus on data readiness, a comprehensive approach should consider the trade-off between data readiness and sufficiency. Sufficient data volumes are essential for meaningful analysis and effective AI model training, making finding the optimal equilibrium between data readiness and quantity a challenge. Additionally, clear interpretability of these readiness metrics is essential for stakeholders to grasp their implications on overall data readiness and their potential impacts on AI model performance. Enhanced visualization and explanation techniques can significantly improve the practical utility of these metrics, facilitating more informed decision-making processes.
@@
Addressing the challenges in the dynamic field of AI and data readiness is important. Specific AI applications often require customized data readiness metrics due to varied data readiness standards, requiring domain-specific expertise for effective navigation. Simultaneously, addressing subjective judgments and human biases, particularly concerning fairness and privacy, adds another layer of complexity. Developing unbiased and ethical metrics that adapt to various data types and applications requires ongoing research and innovation. Furthermore, the ever-evolving nature of real-world datasets requires continuous data readiness assessment throughout an AI system’s life cycle, extending beyond the initial data preparation phase.
@@
Data access and ownership concerns can impede data readiness evaluation, particularly when datasets are restricted due to privacy and ownership issues. These limitations can delay a comprehensive data readiness assessment, requiring collaborative efforts and agreements between data providers and users to navigate these challenges effectively. Furthermore, the ongoing challenge lies in establishing meaningful thresholds that define acceptable data readiness levels. Given the context-dependent nature of data and the diversity of AI use cases, universal and context-independent threshold values are challenging to determine. Clear guidelines regarding data readiness thresholds are essential to ensure consistent and effective data preparation practices. Lastly, developing well-established benchmark datasets and evaluation protocols becomes crucial to compare and evaluate various data readiness metrics. Creating representative benchmarks spanning different industries and data structures can facilitate a more comprehensive comparison of diverse metrics and scoring mechanisms.
Data readiness assessment is critical for all AI applications, including large language models (LLMs)(AI, 2023). Ensuring data completeness, accuracy, and consistency, these metrics will build a robust foundation for LLMs. Correct, unbiased, and relevant data enhance the model’s ability to generate coherent and reliable outputs. Assessing bias in datasets ensures fairness and mitigates skewed predictions. Therefore, applying comprehensive quality metrics is vital to preparing data that effectively supports the latest requirements of LLMs, leading to more accurate and contextually relevant model performance. These requirements highlight the need for more advanced metrics and tools to effectively prepare data to ensure LLMs are trained on high-quality inputs that meet the requirements of modern AI applications.
While having a comprehensive list of metrics in the toolbox is important, deciding the metrics to be used for a particular AI application is essential. When determining the suitable metrics for an AI application, it is crucial to start by defining the objective of the application and being aware of any constraints including data limits or legal considerations. Once the application is defined, one should explore the available metrics and understand each metric and its suitability to achieve the goals. Thet should consider the trade-offs involved and analyze which metrics best align with your objectives. Focusing on metrics that impact the AI application performance and efficiency achieves a balance between simplicity and depth. Ensuring that the AI model’s performance is regularly monitored and that the metrics are flexibly adjusted as it evolves also plays a role in using the correct data. Additionally, reviewing relevant literature can provide valuable insights in selecting the most appropriate metrics. For example, Wagner and Eckhoff (2018) proposed a method for selecting among over $80$ privacy metrics. The selected metrics assess factors such as data sensitivity, trade-offs, and performance expectations, providing meaningful insights and driving the application toward its intended outcomes.
@@ 7. Conclusion
This comprehensive survey underscores the critical role of data preparation toward the effective usage of data by AI applications. We explored the challenges, tools, and metrics associated with data readiness, emphasizing its significance in achieving accurate and dependable AI-driven outcomes. Our study highlights the need for a holistic approach, covering structured and unstructured data, and underscores the importance of incorporating fairness-related metrics to ensure unbiased AI decision-making. We have identified and showcased existing metrics and scoring mechanisms that effectively measure data readiness available in literature by studying more than $140$ publications from reputable journals including ACM, IEEE, and other reputable journals, as well as web articles, spanning the past three decades. By thoroughly exploring these dimensions and metrics, we have summarized numerous data readiness metrics and proposed a taxonomy of them. This will develop a deeper understanding of data readiness for AI applications. As AI advances and data becomes an even more critical asset, staying up to date with the latest research and advancements in data readiness metrics is essential. This survey provides a foundational resource for researchers and practitioners, equipping them with the essential insights needed to navigate the complexities of data preparation for AI effectively and emphasizing that data readiness is not just a preliminary step but an ongoing commitment.